### v1

@@ 1 Introduction
Superfluids and superconductors are macroscopic quantum states of matter that can be described by a single wavefunction with a coherent phase. If two such systems are coupled together, particle currents should oscillate between them, accompanied by oscillations of the relative phase between the two macroscopic wavefunctions. This phenomenon was originally predicted by Josephson in 1962 for superconductors [1] and observed soon after [2, 3]. Since then, the Josephson effect has been studied both theoretically and experimentally in a variety of coupled macroscopic quantum systems, such as superfluid helium reservoirs [4, 5], double well trapped Bose-Einstein condensates [6, 7, 8], dipolar condensates [9, 10, 11, 12], momentum-space condensates [13], magnon condensates [14, 15], polariton condensates [16, 17], supersolids [18, 19, 20], and fermionic superfluids [21, 22, 23, 24].
@@
$\hbar\dot{\phi}$ $=\frac{2Jz}{\sqrt{1-z^{2}}}\cos\phi+UNz,$ (1a)
@@
where $U$ is the boson-boson interaction, $J>0$ is the tunneling energy, and $N$ is the total number of particles [6]. Throughout this work we will consider $U>0$, which corresponds to repulsive interaction. Linearizing around the equilibrium $\phi=z=0$, one finds that small oscillations of $\phi$ and $z$ occur with the frequency
$\omega_{J}=\frac{\sqrt{2J(UN+2J)}}{\hbar}.$ (2)
Qualitatively similar equations govern the dynamics of superconductive Josephson junctions.
In this paper, we address the role of quantum fluctuations and evaluate first-order quantum corrections to the mean-field dynamics described by Eqs. (1) focusing on the relative phase $\phi$, which is directly related to the coherence of the system [25]. We start from the observation that Eqs. (1) are the Euler-Lagrange equations for the action
@@
In the following, we compute analytically the first-order quantum corrections to the dynamics of the phase by first deriving an effective action for $\phi$ integrating out $z$ at the Gaussian level (Section 2) and then computing the corresponding one-loop quantum effective action (Section 3), generalizing a treatment previously presented in Ref. [28]. The range of validity our results, in terms of interaction and number of particles, will be discussed by comparing them with fully quantum numerical results obtained by exact diagonalization of the Bose-Hubbard Hamiltonian (Section 4).
@@ 2 Only-phase effective action
$S^{(2)}[\phi,z]=\int dt\left[-\frac{1}{2m(\phi)}\left(\frac{N\hbar z}{2}\right)^{2}+\frac{N\hbar z}{2}\dot{\phi}-V(\phi)\right]$ (8)
@@
Here $p_{\phi}\equiv N\hbar z/2$ plays the role of conjugate momentum of $\phi$, and the exponential on the right-hand side of Eq. (8) has exactly the form one encounters when computing the phase-space path integral for a quantum particle with coordinate $\phi$ moving in a potential $V(\phi)$ [30]. In this case integrating out $z$ corresponds to the standard passage from the phase-space to the configuration-space path integral. Here, however, the spatial dependence of the mass makes this step nontrivial, for the action $S^{(2)}[\phi,z]$ describes a quantum particle in a curved space with metric $g_{\mu\nu}=m(\phi)\eta_{\mu\nu}$ [31]. Completing the square by shifting the momentum as $\tilde{p}_{\phi}=p_{\phi}-m(\phi)\dot{\phi}$ ($\mathcal{D}\tilde{p}_{\phi}=\mathcal{D}p_{\phi}$), we get
@@
$\mathcal{D}\mu(\phi)\equiv\frac{1}{Z}\prod_{j}\int d\phi_{j}\sqrt{m(\phi_{j})},$ (11)
@@ 3 Quantum effective action
Quantum corrections to the dynamics of $\mathcal{A}[\phi]$ can be computed within the quantum effective action formalism [29, 30], which was introduced many years ago to study the effect of quantum corrections on field theories with spontaneously broken symmetries [32, 33, 34, 35]. The quantum effective action is defined as the Legendre transform $\Gamma[\Phi]=W[J]-\int dt\,J\Phi$} of the generating functional of connected correlation functions, $W[J]=-i\hbar\ln Z[J]=-i\hbar\int\mathcal{D}\mu(\phi)\exp[\frac{i}{\hbar}\int dt\,(\mathcal{A}[\phi]+J\phi)]$, where $\Phi=\delta W[J]/\delta J$ is the quantum average of $\phi$ in the presence of the source $J$. Since $\delta\Gamma[\Phi]/\delta\Phi=-J$, the quantum average $\Phi$ in absence of external sources extremizes $\Gamma[\Phi]$; this is the principle of least action for the full quantum theory. Consequently, the sum of connected diagrams built from the classical action $\mathcal{A}[\phi]+\int dt\,J\phi$ can be obtained from the tree diagrams of the effective action $\Gamma[\Phi]+\int dt\,J\Phi$. In terms of the fluctuations $\eta\equiv\phi-\Phi$, we thus have
@@
The trace in Eq. (21) involves a nonlocal functional of $\Phi(t)$ and cannot be computed exactly. If $\Omega^{2}(\Phi)$ varies slowly in time, however, one can build a derivative expansion around a constant $\Phi$, which yields asymptotically a local expression for $\Gamma_{1}[\Phi]$ [36, 37, 38, 39]. Such adiabatic approximation is valid as long as $|\dot{\Omega}(\Phi)|/\Omega^{2}(\Phi)\ll 1$, and for $\Phi(t)\approx\Phi_{0}\cos(\omega_{J}t)$ with $|\Phi_{0}|\ll 1$ this condition is typically satisfied. One then finds that $\Gamma_{1}[\Phi]$ can be written as the time integral of a Lagrangian, which is a series of terms involving time derivatives of $\Phi(t)$ of increasing order,
@@
With these functions, represented in Fig. 1, Eq. (27) describes the quantum-corrected Josephson dynamics of the relative phase between two coupled condensates. In particular, small oscillations around $\Phi=0$ occur with the frequency
$\Omega_{J}$ $=\sqrt{\frac{V^{\prime\prime}_{\text{eff}}(0)}{m_{\text{eff}}(0)}}=\omega_{J}\sqrt{1-\frac{1}{2N}\frac{UN+6J}{\sqrt{2J(UN+2J)}}}.$ (30)
In the limit $J/UN\to 0$, where $m$ can be approximated to be constant, this reduces to
$\Omega_{J}\simeq\omega_{J}\sqrt{1-\sqrt{\frac{U}{8JN}}},$ (31)
@@ 4 Quantum dynamics of the phase
It is valuable to compare the quantum-corrected Josephson dynamics of Eqs. (27)-(29) with the fully quantum dynamics of the Bose-Hubbard Hamiltonian $\hat{H}$ [Eq. (5)]. In general, quantum evolution depends sensitively on the initial state. To establish a meaningful comparison with a semiclassical dynamics of the collective variables $(\phi,z)$, the initial state should possess well-defined average relative phase and population imbalance. As these are conjugate variables, the natural choice is a state that minimizes the product of their variances. Therefore, we initialize the system at $t=0$ in the $N$-particle atomic coherent state
@@
To extract $\phi_{\text{cond}}(t)$, we first numerically diagonalize $\hat{H}$ to obtain the coefficients $\{c_{k}^{(n)}\}$. We then construct the one-body density matrix and numerically diagonalize it at any given time, identifying the largest-eigenvalue eigenvector. An example of the results we obtain is shown in Fig. 2, where the exact quantum dynamics is compared with the the mean-field and the quantum-corrected dynamics. For the chosen parameter values and initial conditions, $\langle\varrho_{0}(t)\rangle_{\mathrm{time}}/N\simeq 0.94$; both semiclassical dynamics are well described by harmonic oscillations with frequencies $\omega_{J}$ and $\Omega_{J}$, respectively, and the quantum-corrected frequency is $2.1\%$ lower than the classical one. The dynamics obtained from the quantum effective action significantly improves the comparison with the exact result; the agreement is almost perfect until about two oscillation periods, when the amplitude modulation of the oscillations of $\phi_{\text{cond}}(t)$ begins to become evident. This is an expected phenomenon [42] that is not described by the quantum effective action, which nevertheless correctly captures the oscillation frequency.
Figure 3: Range of validity of the quantum-corrected dynamics in terms of $U/J$ and $N$. In region I (yellow) the quantum-corrected dynamics improves the mean field by bringing it closer to the exact dynamics. In region II (green) the quantum-corrected dynamics performs worse than the mean field due to the breakdown of the initial Gaussian integration over $z$. In region III (cyan) the exact dynamics is dominated by strong anharmonicity, and neither the mean field nor the quantum-corrected dynamics provide an accurate description. The transition line between regions I and II fits $\Lambda=UN/2J=10$ (black dashed-dotted line).
@@
We investigated the range of validity of the quantum-corrected dynamics in terms of $U/J$ and $N$, identifying the region in parameter space where it is closer to the exact results than the mean field. This is denoted as region I in Fig. 3. At small $U/J$, i.e. for $\Lambda=UN/2J\lesssim 10$ (region II), the quantum-corrected dynamics loses accuracy due to the breakdown of the assumption underlying the Gaussian integration of $z$, as anticipated in the discussion following Eq. (7). The mean field, however, provides a quite accurate description of this regime, and in the special case $U=0$, $\phi_{\text{cond}}(t)$ follows exactly Eqs. (1) at arbitrary $N$. We remark that this result still depends on the choice of $|\Psi_{\phi_{0},z_{0}}\rangle$ as initial state; if we initialized the system in a Fock state, for example, its quantum evolution could certainly not be described by a single mean-field trajectory, because a Fock state has no well-defined relative phase. However, there are operators whose quantum evolution (as far as only first moments are considered) is described exactly by the mean-field equations for any initial state. This is due to the fact that for $U=0$ the two-site Bose-Hubbard Hamiltonian can be written as $\hat{H}=-2J\hat{S}_{x}$, where $\hat{S}_{x}=\frac{1}{2}(\hat{a}_{L}\hat{a}_{R}+\hat{a}_{R}\hat{a}_{L})$, $\hat{S}_{y}=\frac{1}{2i}(\hat{a}_{L}^{\dagger}\hat{a}_{R}-\hat{a}_{R}^{\dagger}\hat{a}_{L})$, and $\hat{S}_{z}=\frac{1}{2}(\hat{N}_{L}-\hat{N}_{R})$ are spin $N/2$ operators [26]. The unitary operator $e^{-i\hat{H}t/\hbar}=e^{i(2Jt/\hbar)\hat{S}_{x}}$ is therefore a $SU(2)$ rotation around the $\mathbf{e}_{x}$ axis, and for any operator that is a linear combinations of the generators, $\hat{O}=\mathbf{a}\cdot\hat{\mathbf{S}}$, the Heisenberg equation $\partial_{t}\hat{O}(t)=-\frac{i}{\hbar}[\hat{O},\hat{H}]=-2J\mathbf{a}\cdot(\hat{\mathbf{S}}\times\mathbf{e}_{x})$ are also linear. The expectation value $\langle\hat{O}(t)\rangle$ on any initial state thus obeys the same linear equation obtained by replacing the operators by classical variables, $\partial_{t}\langle\hat{O}(t)\rangle=-2J\mathbf{a}\cdot(\langle\hat{\mathbf{S}}\rangle\times\mathbf{e}_{x})$. This is a realization of the Ehrenfest theorem. We finally mention that deep in region II, i.e. for $\Lambda\simeq 1$, and small $N$, the agreement between the mean-field and the exact dynamics can be improved introducing the finite-size correction $U\to U(1-N^{-1})$ in Eqs. (1) [41].
@@
At large $U/J$ (region III), the dynamics of $\phi_{\text{cond}}(t)$ is highly anharmonic and the oscillations’ amplitude is strongly modulated. In this regime, the system’s coherence is reduced (for instance, $\langle\varrho_{0}(t)\rangle_{\mathrm{time}}/N\simeq 0.78$ for $N=50$ and $U/J=5$) and neither the mean-field nor the quantum-corrected dynamics provide an accurate description. The transition between regions II and III is set schematically at $U/J=1$ for $N\leq 25$.

### v2

@@ 1 Introduction
Superfluids and superconductors are macroscopic quantum states of matter that can be described by a single wavefunction. If two such systems are coupled together, particle currents should oscillate between them, accompanied by oscillations of the relative phase between the two macroscopic wavefunctions. This phenomenon was originally predicted by Josephson in 1962 for superconductors [1] and observed soon after [2, 3]. Since then, the Josephson effect has been studied both theoretically and experimentally in a variety of coupled macroscopic quantum systems, such as superfluid helium reservoirs [4, 5], double well trapped Bose-Einstein condensates [6, 7, 8], dipolar condensates [9, 10, 11, 12], momentum-space condensates [13], magnon condensates [14, 15], polariton condensates [16, 17], supersolids [18, 19, 20], and fermionic superfluids [21, 22, 23, 24].
@@
$\hbar\dot{\phi}$ $=\frac{2Jz}{\sqrt{1-z^{2}}}\cos\phi+UNz+\varepsilon,$ (1a)
@@
where $U$ is the boson-boson interaction, $J>0$ is the tunneling energy, $\varepsilon$ is a potential energy difference between the two condensates, and $N$ is the total number of particles [6]. A simplified version of these equations in the regime $|z|\ll 1$, namely $\dot{\phi}=\varepsilon/\hbar$ and $\dot{z}=-(2J/\hbar)\sin\phi$, govern the dynamics of superconductive Josephson junctions, where $\dot{z}$ is proportional to the electric current and $\varepsilon$ to the applied potential difference across the junction [1].
In this paper, we address the role of quantum fluctuations and evaluate first-order quantum corrections to the mean-field dynamics described by Eqs. (1) focusing on the relative phase $\phi$, which is directly related to the coherence of the system [25]. Throughout the work we will consider $U>0$, which corresponds to repulsive interaction, and $\varepsilon=0$. We start from the observation that Eqs. (1) (with $\varepsilon=0$) are the Euler-Lagrange equations for the action
@@
In particular, linearizing around the equilibrium $\phi=z=0$, one finds that small oscillations of $\phi$ and $z$ occur with the frequency
$\omega_{J}=\frac{\sqrt{2J(UN+2J)}}{\hbar}.$ (4)
@@
In the following, we compute analytically the first-order quantum corrections to the dynamics of the phase by first deriving an effective action for $\phi$ integrating out $z$ at the Gaussian level (Section 2) and then computing the corresponding one-loop quantum effective action (Section 3), generalizing a treatment previously presented in Ref. [28]. The range of validity of our results, in terms of interaction and number of particles, will be discussed by comparing them with fully quantum numerical results obtained by exact diagonalization of the Bose-Hubbard Hamiltonian (Section 4).
@@ 2 Only-phase effective action
$S^{(2)}[\phi,z]=\int dt\biggl[-\frac{1}{2m(\phi)}\left(\frac{N\hbar z}{2}\right)^{2}+\frac{N\hbar z}{2}\dot{\phi}-V(\phi)\biggr]$ (8)
@@
Here $p_{\phi}\equiv N\hbar z/2$ plays the role of conjugate momentum of $\phi$, and $\int\mathcal{D}\phi\mathcal{D}z\exp(\frac{i}{\hbar}S^{(2)}[\phi,z])$ has exactly the form one encounters when computing the phase-space path integral for a quantum particle with coordinate $\phi$ moving in a potential $V(\phi)$ [30]. In this case integrating out $z$ corresponds to the standard passage from the phase-space to the configuration-space path integral. Here, however, the spatial dependence of the mass makes this step nontrivial, for the action $S^{(2)}[\phi,z]$ describes a quantum particle in a curved space with metric $g_{\mu\nu}=m(\phi)\eta_{\mu\nu}$ [31]. Completing the square by shifting the momentum as $\tilde{p}_{\phi}=p_{\phi}-m(\phi)\dot{\phi}$ ($\mathcal{D}\tilde{p}_{\phi}=\mathcal{D}p_{\phi}$), we get
@@
$\mathcal{D}\mu(\phi)\equiv\frac{1}{Z}\prod_{j}d\phi_{j}\sqrt{m(\phi_{j})},$ (11)
@@ 3 Quantum effective action
Quantum corrections to the dynamics of $\mathcal{A}[\phi]$ can be computed within the quantum effective action formalism [29, 30], which was introduced many years ago to study quantum effects on field theories with spontaneously broken symmetries [32, 33, 34, 35]. We recall that the quantum effective action is defined as the Legendre transform $\Gamma[\Phi]=W[J]-\int dt\,J\Phi$} of the generating functional of connected correlation functions, $W[J]=-i\hbar\ln Z[J]=-i\hbar\ln\int\mathcal{D}\mu(\phi)\exp[\frac{i}{\hbar}\int dt\,(\mathcal{A}[\phi]+J\phi)]$, where $\Phi=\delta W[J]/\delta J$ is the quantum average of $\phi$ in the presence of the source $J$. Since $\delta\Gamma[\Phi]/\delta\Phi=-J$, the quantum average $\Phi$ in absence of external sources extremizes $\Gamma[\Phi]$; this is the principle of least action for the full quantum theory. Consequently, the sum of connected diagrams built from the classical action $\mathcal{A}[\phi]+\int dt\,J\phi$ can be obtained from the tree diagrams of the effective action $\Gamma[\Phi]+\int dt\,J\Phi$. In terms of the fluctuations $\eta\equiv\phi-\Phi$, we thus have
@@
The trace in Eq. (21) involves a nonlocal functional of $\Phi(t)$ and cannot be computed exactly. If $\Omega^{2}(\Phi)$ varies slowly in time, however, one can build a derivative expansion around a constant $\Phi$, which yields asymptotically a local expression for $\Gamma_{1}[\Phi]$ [36, 37, 38, 39]. Such adiabatic approximation is valid as long as $|\dot{\Omega}(\Phi)|/\Omega^{2}(\Phi)\ll 1$. One then finds that $\Gamma_{1}[\Phi]$ can be written as the time integral of a local Lagrangian, which is a series of terms involving time derivatives of $\Phi(t)$ of increasing order,
@@
With these functions, represented in Fig. 1, Eq. (27) describes the quantum-corrected Josephson dynamics of the relative phase between two coupled condensates. We notice that if $\Phi(t)\sim\Phi_{0}\cos(\omega_{J}t)$, then $|\dot{\Omega}(\Phi)|/\Omega^{2}(\Phi)\sim|\Phi_{0}|$ as order of magnitude, therefore we expect the adiabatic approximation to hold as long as the amplitude of phase oscillations remains much smaller than unity.
Quantum corrections do not change the position of the minimum of the potential, which is still located at $\Phi=0$, where also $m^{\prime}_{\text{eff}}(0)=0$. In particular, small oscillations around $\Phi=0$ are harmonic, with the frequency
$\Omega_{J}$ $=\sqrt{\frac{V^{\prime\prime}_{\text{eff}}(0)}{m_{\text{eff}}(0)}}=\omega_{J}\sqrt{1-\frac{1}{2N}\frac{UN+6J}{\sqrt{2J(UN+2J)}}},$ (30)
where $\omega_{J}$ is the classical frequency (4). In the limit $UN\gg J$, where $m$ can be approximated to be constant, this reduces to
$\Omega_{J}\simeq\frac{\sqrt{2JUN}}{\hbar}\sqrt{1-\sqrt{\frac{U}{8JN}}},$ (31)
@@ 4 Quantum dynamics of the phase
It is valuable to compare the quantum-corrected Josephson dynamics of Eqs. (27)-(29) with the fully quantum dynamics of the Bose-Hubbard Hamiltonian $\hat{H}$ [Eq. (5)]. In general, the quantum evolution depends sensitively on the initial state. To establish a meaningful comparison with a semiclassical dynamics of the collective variables $(\phi,z)$, the initial state should possess well-defined average relative phase and population imbalance. As these are conjugate variables, the natural choice is a state that minimizes the product of their variances. Therefore, we initialize the system at $t=0$ in the $N$-particle atomic coherent state
@@
To extract $\phi_{\text{cond}}(t)$, we first numerically diagonalize $\hat{H}$ to obtain the coefficients $\{c_{k}^{(n)}\}$. Since the dimension of the Hilbert space scales linearly with $N$, this can be easily done up to several thousand particles without the need for particularly sophisticated algorithms. We then construct the one-body density matrix and numerically diagonalize it at any given time, identifying the largest-eigenvalue eigenvector. In practice, we discretize the time interval of interest into $10^{3}$-$10^{5}$ steps. An example of the results we obtain is shown in Fig. 2, where the exact quantum dynamics is compared with the mean-field and the quantum-corrected dynamics (in the figure we use units where $\hbar=J=1$, so that times are also adimensional). For the chosen parameter values and initial conditions, $\langle\varrho_{0}(t)\rangle_{\mathrm{time}}/N\simeq 0.94$, and both semiclassical dynamics are well described by harmonic oscillations with frequencies $\omega_{J}$ and $\Omega_{J}$, respectively, with the quantum-corrected frequency being $2.1\%$ lower than the classical one. The harmonic approximation based on $\Omega_{J}$ remains accurate for oscillation amplitudes up to $0.3$, beyond which the full equation of motion (27) is needed to accurately describe the dynamics. We see that the dynamics obtained from the quantum effective action significantly improves the comparison with the exact result, and the agreement is almost perfect until about two oscillation periods, when the amplitude modulation of the oscillations of $\phi_{\text{cond}}(t)$ starts to become evident. The latter is an expected, genuine many-body quantum effect [42, 43]. In fact, $|\Psi(t)\rangle$ is a superposition of many eigenstates, so expectation values such as $\phi_{\text{cond}}(t)$ contain summations of terms like $e^{-i(E_{n}-E_{m})t/\hbar}$. Because the energy levels of the interacting Hamiltonian are non-equally spaced, a band of closely spaced frequency differences appears and their superposition produces oscillations modulated by a slowly-varying envelope, showing beats, partial collapse and revivals. This mechanism is controlled by $U/J$ and $N$, with stronger interactions and smaller number of particles making the modulations faster and more pronounced (see also the discussion below). This behavior cannot be described by the quantum effective action, which is still a semiclassical description for the collective degrees of freedom of the system; nevertheless, it correctly captures the frequency of fast oscillations over several periods.
@@
We investigated the range of validity of the quantum-corrected dynamics as a function of $U/J$ and $N$, identifying the region in parameter space where it provides a closer agreement with the exact results than the mean-field approximation. This domain, referred to as Region I in Fig. 3, was determined qualitatively by examining the limits in which the quantum-corrected evolution continues to capture the salient features of the exact evolution, notably the oscillation frequency, over at least three periods. Decreasing $U/J$ and increasing $N$ (while remaining within Region I) reduces many-body modulations and allows the maintain quantitative agreement for longer times. Fig. 4 shows representative examples of the dynamics at the boundaries of the validity domain as well as in the parameter regimes where the quantum-corrected results become inaccurate (Regions II and III).
@@
At small $U/J$ (Region II), the quantum-corrected dynamics loses accuracy due to the breakdown of the assumption underlying the Gaussian integration of $z$ (Figs. 4c and 4d), as anticipated in the discussion following Eq. (7). We find that the transition line between Regions I and II fits $\Lambda=10$ for $N\gtrsim 25$. However, the mean field provides a quite accurate description of Region II, and in the special case $U=0$, $\phi_{\text{cond}}(t)$ follows exactly Eqs. (1) at arbitrary $N$. We remark that this result still depends on the choice of $|\Psi_{\phi_{0},z_{0}}\rangle$ as initial state; if we initialized the system in a Fock state, for example, its quantum evolution could certainly not be described by a single mean-field trajectory, because a Fock state has no well-defined relative phase. Instead, there are operators whose quantum evolution (as far as only first moments are considered) is described exactly by the mean-field equations for any initial state. This is due to the fact that for $U=0$ the two-site Bose-Hubbard Hamiltonian can be written as $\hat{H}=-2J\hat{S}_{x}$, where $\hat{S}_{x}=\frac{1}{2}(\hat{a}_{L}\hat{a}_{R}+\hat{a}_{R}\hat{a}_{L})$, $\hat{S}_{y}=\frac{1}{2i}(\hat{a}_{L}^{\dagger}\hat{a}_{R}-\hat{a}_{R}^{\dagger}\hat{a}_{L})$, and $\hat{S}_{z}=\frac{1}{2}(\hat{N}_{L}-\hat{N}_{R})$ are spin $N/2$ operators [26]. The unitary operator $e^{-i\hat{H}t/\hbar}=e^{i(2Jt/\hbar)\hat{S}_{x}}$ is therefore a $SU(2)$ rotation around the $\mathbf{e}_{x}$ axis, and for any operator that is a linear combinations of the generators, $\hat{O}=\mathbf{a}\cdot\hat{\mathbf{S}}$, the Heisenberg equation $\partial_{t}\hat{O}(t)=-\frac{i}{\hbar}[\hat{O},\hat{H}]=-2J\mathbf{a}\cdot(\hat{\mathbf{S}}\times\mathbf{e}_{x})$ are also linear. The expectation value $\langle\hat{O}(t)\rangle$ on any initial state thus obeys the same linear equation obtained by replacing the operators by classical variables, $\partial_{t}\langle\hat{O}(t)\rangle=-2J\mathbf{a}\cdot(\langle\hat{\mathbf{S}}\rangle\times\mathbf{e}_{x})$. This is a realization of the Ehrenfest theorem. Finally, it is worth mentioning that deep in Region II, i.e. for $\Lambda\sim 1$ and small $N$, the agreement between the mean-field and the exact dynamics can be improved by introducing the finite-size correction $U\to U(1-N^{-1})$ in Eqs. (1) [41].
Figure 3: Range of validity of the quantum-corrected dynamics in terms of $U/J$ and $N$. In Region I (yellow) the quantum-corrected dynamics improves the mean field by bringing it closer to the exact dynamics. In Region II (red) the quantum-corrected dynamics performs worse than the mean field due to the breakdown of the initial Gaussian integration over $z$. In Region III (blue) the exact dynamics is dominated by strong anharmonicity, and neither the mean field nor the quantum-corrected dynamics provide an accurate description. The transition line between Regions I and II fits $\Lambda=UN/2J=10$ (black dashed-dotted line).
@@
At large $U/J$ (Region III), the dynamics of $\phi_{\text{cond}}(t)$ is highly anharmonic and the oscillations’ amplitude is strongly modulated on time scales comparable with the inverse of the Josephson frequency. This is a truly quantum regime, where the system’s coherence is reduced due to relatively strong interactions (for instance, $\langle\varrho_{0}(t)\rangle_{\mathrm{time}}/N\simeq 0.78$ for $N=50$ and $U/J=5$) and neither the mean-field nor the quantum-corrected dynamics provide an accurate description (Figs. 4a and 4b). The separation line between Regions II and III can be set schematically at $U/J=1$ for $N\lesssim 25$, which identifies the transition from a weakly interacting and highly coherent system, which is well described by the mean field theory, to a strongly interacting system displaying genuine many-body effects and reduced coherence.
Figure 4: Comparison between the exact dynamics (solid black line), the mean-field dynamics (dashed-dotted blue line), and the quantum-corrected dynamics (dashed red line) of the relative phase, for $N=50$, $J=1.0$, $\phi_{0}=0.1$, and $z_{0}=0$, for several values of $U$ in the regions of parameter space where the quantum-corrected dynamics loses accuracy. (a) $U=5.0$, Region III. (b) $U=2.5$, boundary between Region I and Region III. (c) $U=0.4$, boundary between Region I and Region II. (d) $U=0.05$, Region II. (Units: $\hbar=1$).
Josephson oscillations of the relative phase have been directly observed across several platforms, where interference measurements enable precise tracking of coherent phase dynamics between two weakly coupled modes [7, 16, 17, 23]. Given the current capabilities in controlling tunneling rates, interaction strengths, and atom numbers in bosonic Josephson junctions, it appears experimentally feasible, though technically demanding, to test the theoretical prediction of a reduction in the Josephson oscillation frequency arising from quantum corrections. With dedicated effort aimed at achieving high phase coherence and sub-percent frequency precision, such measurements could provide a direct probe of mesoscopic quantum fluctuations in collective condensate dynamics. However, we emphasize that there already exist indirect evidences of the quantum nature of $\phi$. In superconducting Josephson junctions, this was demonstrated decades ago in pioneering experiments by Clarke and collaborators, who observed macroscopic quantum tunneling and energy-level quantization of the phase [44, 45]. For bosonic Josephson junctions, the condensate fraction (or equivalently, the coherence visibility) constitutes a similarly relevant observable for probing the quantum character of the phase. Within mean-field theory, the ground state of the system always has full condensation, $\varrho_{0}/N=1$, irrespectively of the interaction strength $U$. Indeed, since in mean-field theory the condensate occupation is given by $\varrho_{0}/N=(1+\sqrt{1-z^{2}}\cos\phi)/2$, the minimum of the energy (3) at $\phi=z=0$ yields $\varrho_{0}/N=1$. In contrast, quantum fluctuations introduce a finite zero-point energy that depends on $U$ through the Josephson frequency $\omega_{J}$. The consequence is a reduction of the condensate fraction in the ground state (quantum depletion) [25]. Incorporating the quantum-corrected frequency $\Omega_{J}$ improves the agreement with fully quantum calculations of this depletion. Quantum depletion has already been observed in several ultracold-atom experiments (see e.g. Refs. [46, 47, 48]) and its precise measurement in a double-well geometry would provide an indirect test of the predicted quantum corrections to the Josephson frequency.